Summary statistics for returns on two stocks X and Y are listed below.

Mean Variance
Stock X 2.35% 0.007000
Stock Y 4.53% 0.003000

The covariance of returns on stocks X and Y is 0.002700. Consider a portfolio of 70% stock X and 30% stock Y.

What is the mean of portfolio returns?

Please specify your answer in decimal terms and round your answer to the nearest thousandth (e.g., enter 12.3 percent as 0.123).
Answer #1

mean of portfolio returns =wx*E(x)+wy*E(Y)   (where wx and wy are weights invested in X and Y)

=0.7*2.35+0.3*4.53 =3.004 %